# Multi-task Photonic Reservoir Computing: Wavelength Division Multiplexing for Parallel Computing with a Silicon Microring Resonator

### v1

@@ 5.6 Nonlinearity metric of the MRR-based TDRC
From Eq. 7, we can determine the frequency detuning of the resonance resulting from the nonlinear effects occurring in the MRR, noted $\delta_{\textrm{NL}}(t)$, which is the sum of the frequency detunings due to FCD and the TO effect,:
@@ 6 Results and discussion
Throughout most of this section, the length of both training and testing sets is 2000 elements. The exceptions are Sections 6.5, 6.6 and 6.7.2 where the system addresses different tasks simultaneously. In that case, 10000 symbols are used for training and 10 different subsets of 10000 symbols for testing. The length of the warm-up dataset used before both the training and testing sequences is 250. The value of the regularization parameter ($\Lambda$) used in the training is $1.0\times 10^{-9}$.
@@ 6.2 Multiple optical channels addressing the same task in parallel, M = 4
Here, we first test if the performance of the system shows the same behavior for a $M>1$ number of optical channels over the parameter space of $\overline{P}_{\textrm{i}}$ and $\Delta\omega_{i}$. We address the NARMA-10 task for this test and $M=4$ optical channels for this analysis. For this simulation, $\Delta\phi/2\pi=1/3$, and $\tau_{d}=0.5$ ns, which was optimized in terms of lowest NMSE achieved in simultaneous operation of multiple optical channels (same time delay than in (Giron Castro et al., 2024b)).
@@ 6.2.1 Simultaneous computing of equal instances of the same task
Hence, the total power ($\overline{P}_{\textrm{T}}$) required by the system at the location of the best performance is 5 dBm. This is an order of magnitude higher than in the single-$\lambda$ scenario, but we quadruple the computing capacity of the system for the same task. Another important difference concerning the results of the single-$\lambda$ implementation (as shown in Fig. 2) is that the area where self-pulsing is triggered (affecting the memory of the reservoir) has a significantly narrower extension over the parameter space. It is reduced to a limited range where NMSE $>$ 1.0 in Fig. 3. These differences in the power required to achieve the same level of performance, and in the generation of self-pulsing can be explained by considering the total energy circulating in the cavity.
From the outlined TCMT model, we consider the contribution of the modal amplitude energy of each optical channel $|a_{i}(t)|^{2}$ in the system and compare it with the case where there is a single optical channel. Subsequently, the total modal amplitude energy (corresponding to the single-$\lambda$ implementation) at a given instant is distributed equally over the $M$ optical channels. This is ($|a_{i}(t)|^{2}=|a(t)|^{2}/M$. Then, the energy contribution from each optical channel to the change rate of $\Delta N(t)$ in the WDM scenario at such instant is $1/M^{2}$ of the value of the single-$\lambda$ case as determined by $(|a(t)|^{2}/M)^{2}=|a(t)|^{4}/M^{2}$.
We can infer from Eq. 8 that a lower change rate of $\Delta N(t)$ also affects the FCA generation and the conversion of absorbed power into heat and in turn, decreases the change rate of $\Delta T(t)$. From, Eq. 22, we see that this reduction in the generation rate of both $\Delta N(t)$ and $\Delta T(t)$ causes an overall lower nonlinear detuning for the WDM MRR-based TDRC. In summary, more power is required by the WDM-based system to achieve the same level of nonlinear dynamics as the single-channel system.
Consequently, triggering self-pulsing (Region C in Fig. 2) in the WDM-based TDRC would require additional power than if we used a single optical channel. This difference in energy appears to make the system avoid self-pulsing. As self-pulsing is detrimental to the memory of the RC, avoiding it is beneficial for the WDM MRR-based TDRC. Nevertheless, this also means that more power is required to achieve a moderate level of nonlinear dynamics where the system performs the best (Region B).
@@ 6.2.2 Simultaneous computing of different instances of the same task
There is a decrease in performance with respect to the case where we used the same set of seeds to generate the modulating signals of every optical channel. This is to be expected if we consider that the carrier population interacts with all the channels through the cavity nonlinear dynamics. If the modulated optical signals are equal, then the short memory provided by the carrier population will contain information that favors equally the computing processing in every channel. This effect appears to also compensate for the impact on the system of the resulting phase shift in the feedback mechanism (further studied in Section 6.7). Therefore, the inter-coupling between the nodes of each optical channel benefits the all-around performance of the system when using the same sets of seeds. This of course, is not the case when using different sets of seeds.
We can see that the region of best performance over the parameter space becomes narrower for this case (Fig. 4) and the difference of NMSE between the channels, particularly between channels 0 and 1 is increased (see Supplementary material). However, we still obtain very good performance for the NARMA-10 task in every instance. The minimum NMSE in every multiplexed channel is different, but the location over the parameter space remains the same for all. In fact, it is the same location as in the previously studied scenario $(\overline{P}_{\textrm{i}}\approx-1.02$ dBm and $\Delta\omega_{i}/2\pi=\nolinebreak 35\textrm{GHz})$.
@@ 6.3 Performance vs number of channels for the NARMA-10 task
This issue can be addressed by modulating the respective data of channels $\omega_{6}$ and $\omega_{7}$ to successive resonances of the MRR ($i=\{-5,6\}$). We simulated the system with this change of frequencies of the channels and obtained the result shown in Fig. 6 for $\overline{P}_{\textrm{T}}=0$ dBm.
@@
Now all the channels exhibit similar performance. As studied in Section 6.7, a potential alternative to address this problem is optimizing the total phase shift in the delay waveguide.
@@ 6.4 Memory capacity of the WDM MRR-based photonic RC, M = 4
Another method to validate the suitability of each optical channel for computing purposes without depending on a particular task is to calculate the memory capacity, as previously defined in Section 5. In this case, the target of RC is to reconstruct the input that was used previously to generate the NARMA-10 sequence, for $K_{\textrm{max}}=50$ and $H=3$. The rest of the parameters used in this simulation are the same as in Section 6.2 for $M=4$.
The results are visible in Fig. 7. It is clear that the $MC$ achieved by all optical channels is very similar over the parameter space of $\overline{P}_{\textrm{in}}$ and $\Delta\omega_{i}$. As the $MC$ is independent of a particular task, we can infer that for $M$ up to 4, the memory capabilities of each optical channel in this scheme are approximately the same.
@@
When considering different tasks in each optical channel, a certain degree of freedom is available. The $MC$ varies over the parameter space and this could allow adjusting each optical channel to the memory that its respective task demands. However, this flexibility is limited as the nonlinear dynamics of the MRR are intertwined to each optical channel, e.g. through (8)-(9); hence, a change in $\overline{P}_{\textrm{in}}$ or $\Delta\omega_{i}$ in one optical channel to optimize its respective $MC$, would affect the performance of the others.
@@ 6.5 Multiple optical channels addressing different tasks in parallel, M = 4
The initial part of this section focused on the capabilities of the studied RC scheme to reproduce similar performance when addressing simultaneously the same task in different optical channels. Now we focus on addressing different tasks on each of the optical channels, following the setup of the tasks shown in Fig. 1. For this purpose, we set the value $M=4$ as a proof of concept, where the angular frequency of the $i^{\textrm{th}}$ modulated optical channel follows the expression $\omega_{i}=\omega_{0}+i\cdot\textrm{FSR}_{\omega}$ for $i=\{-1,0,1,2\}$. To simplify the simulation process, the system uses the same data lengths for the three tasks, 10000 symbols for training and 10 different subsets of 10000 symbols for testing. $\Delta\phi/2\pi$ is set to 2/3.
@@
The result of each multiplexed task is shown in Fig. 8. Similar to our initial insights in (Giron Castro et al., 2024c), when using the same value of $\overline{P}_{i}$ and $\Delta\omega_{i}$ for every channel, the best performance of each task differs considerably in their respective location in the parameter space. Therefore, no simultaneous operation with optimal performance of the system is achievable in the previously simulated conditions.
Consequently, we select the power and detuning values leading to the best performance for each individual task from Fig. 8 and evaluate the simultaneous performance under those conditions. As concluded in (Giron Castro et al., 2024c), this comes at the cost of a performance penalty in each task, consequently setting up a trade-off between performance and parallelization for the WDM MRR-based TDRC scheme. The focus of this work is on the parallelization of the tasks and not on the absolute best performance. Nonetheless, we still compare the results obtained with those of previous works on photonic TDRC. We achieve simultaneous operation of the tasks by adjusting the values of the input power and angular frequency to those listed in Table 2.
@@
We list the results regarding the best performance of each task and compare them with those of previous studies on photonic TDRC (see Supplementary material). We point out that the results of the previous studies are for photonic reservoirs that address a single task at a time. Indeed, there is a small performance penalty even when compared to the single-$\lambda$ MRR-based TDRC. Although outside of the scope of this study, this penalty could be reduced with a full 8-D optimization of the values of $\overline{P}_{i}$ and $\Delta\omega_{i}$. Another relevant aspect of this implementation is the higher processing rate of data concerning the other works in the literature. It is also important to highlight that the proposed system requires further optimization of the frequency detuning and input power, just when running different tasks. If we use this system to address multiple times the same task in parallel, as in Section 6.2, we can likely avoid these additional complications.
@@ 6.6 Nonlinearity of each optical channel for different tasks
As anticipated, the nonlinear detuning of the optical channels follows a very similar pattern. The thermal effect becomes dominant as the power increases, and subsequently, the area of the parameter space with the highest value of $\sigma(\delta_{\textrm{NL}}(t))$ red-shifts. Henceforth, an asymmetry in both the nonlinear detuning and the performance of the tasks is formed over the parameter space of $\overline{P}_{\textrm{i}}$ and $\Delta\omega_{i}/2\pi$, as earlier discussed for the NARMA-10 task in Section 6.3. Comparing the nonlinear detuning to that of the single-$\lambda$ scheme ((Giron Castro et al., 2024a), we see that the maximum nonlinear detuning achieved for the WDM MRR-based TDRC is around five orders of magnitude lower when compared over the same range of input power and $\Delta\omega$, which as discussed before, is a consequence of the different distribution of the energy in the MRR cavity.
A more in-depth analysis of the nonlinear detuning is necessary to verify if self-pulsing is triggered when adjusting independently the input power and frequency detuning of each channel. This also could potentially provide further generalization on the locations over the parameter space that are more favorable for a particular task.
Ultimately, as demonstrated by the results of Figs. 7, 8 and 9, the difference in performance between tasks for a given value of $\overline{P}_{\textrm{i}}$ and $\Delta\omega_{i}/2\pi$ does not seem related to a difference in memory or nonlinearity properties of each channel, as the behavior among channels of both properties over the parameter space is roughly the same. The difference most likely relies only on the computing requirements of each task.
@@ 6.7.1 Impact of $\Delta\phi$ when addressing multiple instances of the same task
In Fig. 5 (d) and Fig. 6, we observed how, for some optical channels the performance over the parameter space differed from that of other optical channels, when addressing the same task. We have attributed this behavior so far to the wavelength dependence of the phase shift in the delay waveguide. To confirm this hypothesis, we simulate the system when varying the value of $\Delta\phi$ for $M=10$, $\overline{P}_{\textrm{T}}=0$ dBm, addressing the NARMA-10 task in all the channels. The results are shown in Fig. 10. We observe how each value of $\Delta\phi$ changes the optical channels that perform noticeably different as a function of $\Delta\omega_{i}$. For $\Delta\phi/2\pi=0$ (Fig. 10 (a)), the affected channels are $\omega_{2}$ and $\omega_{7}$. For $\Delta\phi/2\pi=1/4$ (Fig. 10 (b)), the affected channel is $\omega_{6}$. In the case of $\Delta\phi/2\pi=1/2$ (Fig. 10 (c)), it is $\omega_{0}$. Finally, for $\Delta\phi/2\pi=2/3$ (Fig. 10 (d)) it is $\omega_{1}$ that performs differently.
@@
Interestingly, the impact of $\phi_{i}$ on a particular channel does not necessarily mean a performance penalty. In fact, we can see that in Fig. 10 (b), the optical channel that is affected performs better than the rest, and for a considerably larger range of $\Delta\omega_{i}$. The results correspond to a limited set of $\Delta\phi$, but they already convey the importance of the total phase shift experienced by each channel ($\phi_{i}$) in determining its performance. As also inferred by the next analysis, the control of the phase shift appears to be a key component in improving the overall performance of the WDM MRR-based TDRC.

### v2

@@ 5.6 Nonlinearity metric of the MRR-based TDRC
From Eq. 7, we can determine the frequency detuning of the resonance resulting from the nonlinear effects occurring in the MRR, noted $\delta_{\textrm{NL}}(t)$, which is the sum of the frequency detunings due to FCD and the TO effect, given by Eq. 26:
@@ 6 Results and discussion
Throughout most of this section, the length of both training and testing sets is 2000 elements. The exceptions are Sections 6.5, 6.6 and 6.7.2 where the system addresses different tasks simultaneously. In that case, 10000 symbols are used for training and 10 different subsets of 10000 symbols for testing. The length of the warm-up dataset used before both the training and testing sequences is 250.
@@ 6.2 Multiple optical channels addressing the same task in parallel, M = 4
Here, we first test if the performance of the system shows the same behavior for a $M>1$ number of optical channels over the parameter space of $\overline{P}_{\textrm{i}}$ and $\Delta\omega_{i}$. We address the NARMA-10 task for this test and $M=4$ optical channels for this analysis. For this simulation, $\Delta\phi/2\pi=1/3$, and $\tau_{d}=0.5$ ns, which was optimized in terms of the lowest NMSE achieved in simultaneous operation of multiple optical channels (same time delay than in (Giron Castro et al., 2024b)). Every task uses the same value of bias, $\beta=8.0$.
@@ 6.2.1 Simultaneous computing of equal instances of the same task
Hence, the total power ($\overline{P}_{\textrm{T}}$) required by the system at the location of the best performance is 5 dBm. This is an order of magnitude higher than in the single-$\lambda$ scenario, but we quadruple the computing capacity of the system for the same task. Another important difference concerning the results of the single-$\lambda$ implementation (as shown in Fig. 2) is that the area where self-pulsing is triggered (affecting the memory of the reservoir) has a significantly narrower extension over the parameter space. It is reduced to a limited range where NMSE $>$ 1.0 in Fig. 3. The existence of this limited range of self-pulsing in the negative detuning region, which corresponds to a red shift of the resonance, hints at dominance of the thermo-optic effect when the total power is increased. The shift of region C to negative detuning when the thermo-optic effect is dominant was also observed in our previous study of a single-channel MRR-based TDRC (Giron Castro et al., 2024b).
These differences in the power required to achieve the same level of performance, and in the generation of self-pulsing can be explained by considering the total energy circulating in the cavity. From the outlined TCMT model, we consider the contribution of the modal amplitude energy of each optical channel $|a_{i}(t)|^{2}$ in the system and compare it with the case where there is a single optical channel. Subsequently, the total modal amplitude energy (corresponding to the single-$\lambda$ implementation) at a given instant is distributed equally over the $M$ optical channels. This is ($|a_{i}(t)|^{2}=|a(t)|^{2}/M$. Then, the energy contribution from each optical channel to the change rate of $\Delta N(t)$ in the WDM scenario at such instant is $1/M^{2}$ of the value of the single-$\lambda$ case as determined by $(|a(t)|^{2}/M)^{2}=|a(t)|^{4}/M^{2}$.
We can infer from Eq. 8 that a lower change rate of $\Delta N(t)$ also affects the FCA generation and the conversion of absorbed power into heat and in turn, decreases the change rate of $\Delta T(t)$. From Eq. 26, we see that this reduction in the generation rate of both $\Delta N(t)$ and $\Delta T(t)$ causes an overall lower nonlinear detuning for the WDM MRR-based TDRC. In summary, more power is required by the WDM-based system to achieve the same level of nonlinear dynamics as the single-channel system. Consequently, triggering self-pulsing (Region C in Fig. 2) in the WDM-based TDRC would require additional power than if we used a single optical channel. This difference in energy appears to reduce the area over the parameter space where self-pulsing occurs. This is beneficial for the system as self-pulsing is detrimental to the memory of the RC as discussed later in Section 6.4. Nevertheless, this also means that more power is required to achieve a moderate level of nonlinear dynamics where the system performs the best (Region B).
@@ 6.2.2 Simultaneous computing of different instances of the same task
There is a decrease in performance with respect to the case where we used the same set of seeds to generate the modulating signals of every optical channel. This is to be expected if we consider that the carrier population interacts with all the channels through the cavity nonlinear dynamics. If the modulated optical signals are equal, then the short memory provided by the carrier population will contain information that favors equally the computing processing in every channel. This effect appears to also compensate for the impact on the system of the resulting phase shift in the feedback mechanism (further studied in Section 6.7). Henceforth, the inter-coupling between the nodes of each optical channel benefits the all-around performance of the system when using the same sets of seeds.
This benefit is not present when using different sets of seeds. However, in that scenario, it does not translate to a performance penalty either. The imprinting of unwanted information between the tasks is averaged out in the system due to the difference in speed between the input and the lifetime of the nonlinear effects. This is also the case when the system addresses different tasks in parallel as shown later in Section 6.5.
We can see that the region of the best performance over the parameter space becomes narrower for this case (Fig. 4) and the difference of NMSE between the channels, particularly between channels 0 and 1, is increased (see Supplementary material). However, we still obtain very good performance for the NARMA-10 task in every instance. The minimum NMSE in every multiplexed channel is different, but the location over the parameter space remains the same for all. In fact, it is the same location as in the previously studied scenario $(\overline{P}_{\textrm{i}}\approx-1.02$ dBm and $\Delta\omega_{i}/2\pi=\nolinebreak 35\textrm{GHz})$.
@@ 6.3 Performance vs number of channels for the NARMA-10 task
This issue can be addressed by modulating the respective data of channels $\omega_{6}$ and $\omega_{7}$ to successive resonances of the MRR ($i=\{-5,6\}$). We simulated the system with this change of frequencies of the channels and obtained the result shown in Fig. 6 for $\overline{P}_{\textrm{T}}=0$ dBm. Now all the channels exhibit similar performance. As studied in Section 6.7, a potential alternative to address this problem is optimizing the total phase shift in the delay waveguide.
@@ 6.4 Memory capacity of the WDM MRR-based photonic RC, M = 4
Another method to validate the suitability of each optical channel for computing purposes without depending on a particular task is to calculate the memory capacity, as previously defined in Section 5. In this case, the target of RC is to reconstruct the input that was used previously to generate the NARMA-10 sequence, for $K_{\textrm{max}}=50$ and $H=3$. The rest of the parameters used in this simulation are the same as in Section 6.2 for $M=4$. The results are visible in Fig. 7. It is clear that the $MC$ achieved by all optical channels is very similar over the parameter space of $\overline{P}_{\textrm{in}}$ and $\Delta\omega_{i}$. As the $MC$ is independent of a particular task, we can infer that for $M$ up to 4, the memory capabilities of each optical channel in this scheme are approximately the same.
@@
When considering different tasks in each optical channel, a certain degree of freedom is available. The $MC$ varies over the parameter space and this could allow adjusting each optical channel to the memory that its respective task demands. However, this flexibility is limited as the nonlinear dynamics of the MRR are intertwined to each optical channel, e.g. through Eqs. 8 and 9; hence, a change in $\overline{P}_{\textrm{in}}$ or $\Delta\omega_{i}$ in one optical channel to optimize its respective $MC$, would affect the performance of the others.
@@ 6.5 Multiple optical channels addressing different tasks in parallel, M = 4
The initial part of this section focused on the capabilities of the studied RC scheme to reproduce similar performance when addressing simultaneously the same task in different optical channels. Now we focus on addressing different tasks on each of the optical channels, following the setup of the tasks shown in Fig. 1. For this purpose, we set the value $M=4$ as a proof of concept, where the angular frequency of the $i^{\textrm{th}}$ modulated optical channel follows the expression $\omega_{i}=\omega_{0}+i\cdot\textrm{FSR}_{\omega}$ for $i=\{0,1,2,3\}$. To simplify the simulation process, the system uses the same data lengths for the three tasks, 20000 symbols for training and 10 different subsets of 10000 symbols for testing. $\Delta\phi/2\pi$ is set to 2/3. The corresponding biases per task are $\beta_{i}=\{8.0,4.0,11.0,10.0\}$. The result of each multiplexed task is shown in Fig. 8.
@@
Similar to our initial insights in (Giron Castro et al., 2024), when using the same value of $\overline{P}_{i}$ and $\Delta\omega_{i}$ for every channel, the best performance of each task differs considerably in their respective location in the parameter space. Therefore, no simultaneous operation with optimal performance of the system is achievable in the previously simulated conditions. Consequently, we simulated the system under different values of power and detuning values across the individual tasks. As a basis, we used the regions that exhibit the best performance over the parameter space in Fig. 8. In Table 2, we list the selected values where simultaneous performance with good performance in each task was achieved. As concluded in (Giron Castro et al., 2024), this can come at the cost of a slight performance penalty, consequently setting up a trade-off between performance and parallelization for the WDM MRR-based TDRC scheme. The focus of this work is on the parallelization of the tasks and not on the absolute best performance. Nonetheless, we still compare the results obtained with those of previous works on photonic TDRC. We achieve simultaneous operation of the tasks by adjusting the values of the input power and angular frequency to those listed in Table 2.
@@
We list the results regarding the best performance of each task and compare them with those of previous studies on photonic TDRC (see Supplementary material). We point out that the results of the previous studies are for photonic reservoirs that address a single task at a time. Indeed, there is a small performance penalty even when compared to the single-$\lambda$ MRR-based TDRC. Although outside of the scope of this study, this penalty could be reduced with a full 8-D optimization of the values of $\overline{P}_{i}$ and $\Delta\omega_{i}$. Another relevant aspect of this implementation is the higher processing rate of data concerning the other works in the literature. We also highlight that the proposed system requires further optimization of the frequency detuning and input power, only when running different tasks. If we use this system to address multiple times the same task in parallel, as in Section 6.2, we can likely avoid these additional complications.
As in Section 6.2, there is an interaction between the tasks because they are under the influence of the same nonlinear dynamics. A change in $\overline{P}_{\textrm{i}}$ or $\Delta\omega_{i}/2\pi$ of one of the channels affects the performance of the others by changing the temporal dynamics of the cavity. However, the output layer of each task is trained using the higher dimensional states of the RC in each channel that result from the nonlinear transformation. Hence, the training addressing each task is bounded by the specific number of tasks present and the parameters of the optical channels used in the setup. The reservoir must be retrained if this configuration changes. This is not different from any other type of RC scheme. It is important to highlight that the training and testing datasets are uncorrelated and separated by a warm-up sequence. Consequently, any relation between the tasks learned during the training phase is randomized and does not impact the presented results, which are based on the testing sequence. Therefore, the good testing performance proves the effectiveness of multi-task processing through WDM.
@@ 6.6 Nonlinearity of each optical channel for different tasks
As anticipated, the nonlinear detuning of the optical channels follows a very similar pattern. The thermal effect becomes dominant as the power increases, and subsequently, the area of the parameter space with the highest value of $\sigma(\delta_{\textrm{NL}}(t))$ red-shifts. Henceforth, an asymmetry in both the nonlinear detuning and the performance of the tasks is formed over the parameter space of $\overline{P}_{\textrm{i}}$ and $\Delta\omega_{i}/2\pi$, as earlier discussed for the NARMA-10 task in Section 6.3. Comparing the nonlinear detuning to that of the single-$\lambda$ scheme ((Giron Castro et al., 2024a), we see that the maximum nonlinear detuning achieved for the WDM MRR-based TDRC is around five orders of magnitude lower when compared over the same range of $\overline{P}_{\textrm{i}}$ and $\Delta\omega_{i}$, which as discussed before, is a consequence of the different distribution of the energy in the MRR cavity.
A more in-depth analysis of the nonlinear detuning is necessary to verify if self-pulsing is triggered when adjusting independently the input power and frequency detuning of each channel. This could provide further generalization on the locations over the parameter space that are more favorable for a particular task. Ultimately, as demonstrated by the results of Figs. 7, 8 and 9, the difference in performance between tasks for a given value of $\overline{P}_{\textrm{i}}$ and $\Delta\omega_{i}/2\pi$ does not seem related to a difference in memory or nonlinearity properties of each channel, as the behavior among channels of both properties over the parameter space is roughly the same. The difference likely relies only on the computing requirements of each task.
@@ 6.7.1 Impact of $\Delta\phi$ when addressing multiple instances of the same task
In Fig. 5 (d) and Fig. 6, we observed how, for some optical channels the performance over the parameter space differed from that of other optical channels, when addressing the same task. We have attributed this behavior so far to the wavelength dependence of the phase shift in the delay waveguide. To confirm this hypothesis, we simulate the system when varying the value of $\Delta\phi$ for $M=10$, $\overline{P}_{\textrm{T}}=0$ dBm, addressing the NARMA-10 task in all the channels. The results are shown in Fig. 10.
@@
We observe how each value of $\Delta\phi$ changes the optical channels that perform noticeably different as a function of $\Delta\omega_{i}$. For $\Delta\phi/2\pi=0$ (Fig. 10 (a)), the affected channels are $\omega_{2}$ and $\omega_{7}$. For $\Delta\phi/2\pi=1/4$ (Fig. 10 (b)), the affected channel is $\omega_{6}$. In the case of $\Delta\phi/2\pi=1/2$ (Fig. 10 (c)), it is $\omega_{0}$. Finally, for $\Delta\phi/2\pi=2/3$ (Fig. 10 (d)) it is $\omega_{1}$ that performs differently. Interestingly, the impact of $\phi_{i}$ on a particular channel does not necessarily mean a performance penalty. In fact, we can see that in Fig. 10 (b), the optical channel that is affected performs better than the rest, and for a considerably larger range of $\Delta\omega_{i}$. The results correspond to a limited set of $\Delta\phi$, but they already convey the importance of the total phase shift experienced by each channel ($\phi_{i}$) in determining its performance. As also inferred by the next analysis, the control of the phase shift appears to be a key component in improving the overall performance of the WDM MRR-based TDRC.